OpenStudy (anonymous):

6) If sec theta+tan theta = P. PT sin theta=P^2-1/P^2+1

myininaya (myininaya):

we want to show the second part?

OpenStudy (anonymous):

yea!

myininaya (myininaya):

what is T?

OpenStudy (anonymous):

PT= PROVE THAT

myininaya (myininaya):

so we get to suppose \[\sec(\theta)+\tan(\theta)=P\] and we want to show \[PT \sin(\theta)=\frac{P^2-1}{P^2+1}\]?

myininaya (myininaya):

oh ok

OpenStudy (anonymous):

and guys can you answer my prev.Questions too?! please..

myininaya (myininaya):

\[\sec(\theta)+\tan(\theta)=\frac{1}{\cos(\theta)}+\frac{\sin(\theta)}{\cos(\theta)}=\frac{1+\sin(\theta)}{\cos(\theta)}=P\] thinking....

myininaya (myininaya):

\[\frac{1+\sin(\theta)}{\cos(\theta)}*\frac{1-\sin(\theta)}{1-\sin(\theta)}=\frac{1-\sin^2(\theta)}{\cos(\theta)(1-\sin(\theta))}\] \[=\frac{\cos^2(\theta)}{\cos(\theta)(1-\sin(\theta))}=\frac{\cos(\theta)}{1-\sin(\theta)}=P\] hmmm....

OpenStudy (anonymous):

solved?

myininaya (myininaya):

no lol

myininaya (myininaya):

i havent solved it yet

myininaya (myininaya):

maybe i should try this on paper first

OpenStudy (anonymous):

what bout my prev.sums? can you solve it?

myininaya (myininaya):

let me try this one first

OpenStudy (anonymous):

sure! :)

myininaya (myininaya):

if \[\sec(\theta)+\tan(\theta)=P\] then \[(\sec(\theta)+\tan(\theta))(\sec(\theta)-\tan(\theta))=P(\sec(\theta)-\tan(\theta))\] but \[(\sec(\theta)+\tan(\theta))(\sec(\theta)-\tan(\theta))=\sec^2(\theta)-\tan^2(\theta)=1\] so we have \[1=P(\sec(\theta)-\tan(\theta))\] => \[\sec(\theta)-\tan(\theta)=\frac{1}{P}\] lets see if we can use this

myininaya (myininaya):

lets see we have that \[\sec(\theta)+\tan(\theta)=\frac{1+\sin(\theta)}{\cos(\theta)}\] and \[\sec(\theta)-\tan(\theta)=\frac{1-\sin(\theta)}{\cos(\theta)}\] adding these together we get \[\frac{1+\sin(\theta)}{\cos(\theta)}+\frac{1-\sin(\theta)}{\cos(\theta)}=P+\frac{1}{P}\] \[\frac{2}{\cos(\theta)}=\frac{P^2+1}{P}\] \[\frac{2P}{P^2+1}=\cos(\theta)\]

myininaya (myininaya):

but if we imagine we have a right triangle and we draw theta somewhere then cos tells adj/hyp=2p/(P^2+1) we can find the opposite of side of theta opp^2+adj^2=hyp^2 \[opp^2=hyp^2-adj^2=(P^2+1)^2-(2P)^2=P^4+2P^2+1-2^2P^2\] so we have \[opp=\sqrt{P^4-2P^2+1}\] but sin=opp/hyp

myininaya (myininaya):

\[\sin(\theta)=\frac{\sqrt{P^4-2P^2+1}}{P^2+1}=\frac{\sqrt{(P^2-1)^2}}{P^2+1}\]

myininaya (myininaya):

\[\sin(\theta)=\frac{P^2-1}{P^2+1}\]

myininaya (myininaya):

any questions?

OpenStudy (anonymous):

yea..can you solve the other problems?

myininaya (myininaya):

zarkon?

OpenStudy (anonymous):

what?

myininaya (myininaya):

i was looking for someone

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

and can you temme one thing..can you temme the stepin correct order so that i can copy it in my rough boook! ;)

myininaya (myininaya):

hey zarkon i had to think on it isn't that crazy?

OpenStudy (zarkon):

yes

myininaya (myininaya):

the problem

OpenStudy (zarkon):

trying to think if there is an easier way

myininaya (myininaya):

i have no clue it was hard for me to see that way

myininaya (myininaya):

what is temme?

myininaya (myininaya):

i don't know that word

myininaya (myininaya):

use small words please lol

myininaya (myininaya):

well i mean i guess its small but i don't know it

myininaya (myininaya):

tell me?

myininaya (myininaya):

temme=tell me?

myininaya (myininaya):

i thought i showed u the steps zarkon thinks there may be an easier way though

myininaya (myininaya):

zarkon i did this for you :)

myininaya (myininaya):

lol

OpenStudy (zarkon):

;)

OpenStudy (zarkon):

how about this...

myininaya (myininaya):

and for abichu dont feel left out

OpenStudy (zarkon):

get to \[\frac{1+sin(\theta)}{\cos(\theta)}=p\] ...

OpenStudy (anonymous):

i am soo weak in math..that's why am seeking help from you guys! if you dont mind help me! ;)

myininaya (myininaya):

we are helping

OpenStudy (zarkon):

\[1+\sin(\theta)=p\cos(\theta)\]

myininaya (myininaya):

i showed you one way and zarkon is gonna show you another it appears

OpenStudy (zarkon):

\[1+\sin(\theta)=p\sqrt{1-\sin^2(\theta)}\]

OpenStudy (zarkon):

square both sides...

myininaya (myininaya):

omg i see where you are going

OpenStudy (zarkon):

\[1+2\sin(\theta)+\sin^2(\theta)=p^2(1-\sin^2(\theta))\]

myininaya (myininaya):

you going to write is as a quadratic

myininaya (myininaya):

and solve for sin(theta)

OpenStudy (anonymous):

can you tel me from the 1st step if you dot mind?!

OpenStudy (zarkon):

move stuff to one side \[1-p^2+2\sin(\theta)+(1+p^2)\sin^2(\theta)=0\]

myininaya (myininaya):

brillant!

OpenStudy (zarkon):

\[1-p^2+2x+(1+p^2)x^2=0\]

OpenStudy (zarkon):

solve this quadratic ;)

OpenStudy (zarkon):

fun stuff :)

myininaya (myininaya):

yes

myininaya (myininaya):

you are a showoff lol

OpenStudy (zarkon):

lol...now abichu has two ways to do it

OpenStudy (anonymous):

i donno how to solve it quadratic..can you doo it n give?

myininaya (myininaya):

\[x=\frac{-2 \pm \sqrt{(2)^2-4(1+p^2)(1-p^2)}}{2(1+p^2)}\]